FTIR AND DFT STUDY OF THE VIBRATIONAL SPECTRUM OF SiC_{5} TRAPPED IN SOLID Ar

\underline{T. H. LE}, and W. R. M. GRAHAM, Department of Physics and Astronomy, Texas Christian University, Fort Worth, TX 76129.

This paper follows preliminary results on the SiC_{5} molecule presented earlier. An absorption at 936.9 ± 0.2 cm^{−1} has been identified as the ν_{4}(σ) Si-C stretching fundamental of linear SiC_{5}. Its assignment has been confirmed by the close agreement between DFT predicted and observed ^{13}C isotopic shifts. DFT-B3LYP calculations predict that two other C-C fundamentals, ν_{1}(σ) and ν_{2}(σ) of SiC_{5} should have intensities 10 to 15 times stronger than ν_{4}(σ), but would appear in the C-C stretching frequency regions. Refinements of the Si-C laser ablation techniques have helped to reveal the isotopic^{13}C shifts for two additional absorptions at 2045.0 and 1992.9 cm^{−1}. Comparison of their observed isotopic shifts with the predictions of DFT calculations at the MPW1PW91/6-311+G(3fd) level have led to their assignment to the ν_{1}(σ) and ν_{2}(σ) fundamentals of SiC_{5}.